OpenStudy (anonymous):

How do I solve this problem? The volumes of two cylindrical cans of the same shape vary directly as the cubes of their radii. If a can with a six-inch radius holds 1½ pints, how many gallons will a similar can with a 24-inch radius hold?

OpenStudy (dan815):

how many pints is a gallon

OpenStudy (anonymous):

I don't know, I'll have to google it.

OpenStudy (dan815):

okay do that

OpenStudy (dan815):

now the volumes of these 2 objects vary as a function of the cubes of their radii

OpenStudy (anonymous):

8

OpenStudy (dan815):

lets suppose Volume of Can A = k*r^3 volume of can B = k*R^3 where R= 24 and r=6 we want to see how many of Can A fit in can B so we know hoiw many pints or gallons go in there

OpenStudy (anonymous):

Oh, I just feel so stressed and it makes difficult for me to comprehend.

OpenStudy (dan815):

okay dont worry we will take this slow

OpenStudy (anonymous):

So, what's the very first step in this problem?

OpenStudy (dan815):

How do I solve this problem? The volumes of two cylindrical cans of the same shape vary directly as the cubes of their radii.

OpenStudy (dan815):

The volumes of two cylindrical cans of the same shape vary directly as the cubes of their radii.

OpenStudy (dan815):

read this sentence

OpenStudy (anonymous):

It says you've been typing for a while. Is that an error?

OpenStudy (dan815):

let Vol_A be the volume of can with radius r = 6 inches Vol_A=k*r^3 Vol_A is varying with respect to r^3 so some constant k times r^3

OpenStudy (anonymous):

Is there a significance to your typing of the _ symbol?

OpenStudy (dan815):

no =.= its just a label for the variable

OpenStudy (dan815):

let Vol_B be the volume of can with radius R=24 inches Vol_B=k*R^3 Vol_B is varying with respect to R^3 so some constant k(notice same constant) times R^3

OpenStudy (dan815):

now what we want to know is how many times bigger volume B is than volume A

OpenStudy (dan815):

then we will know how many of Vol_A will fit in volume_B

OpenStudy (dan815):